ADNAN MUHAMMED

Grade 12,

1]If A = B+C and the magnitude of A, B and C are 5,4,3 unit, the angle between A and B is : (a)cos -1 3/5 (b) cos -1 4/5 (c)π/2 (d) sin -1 3/5 2] Unit vector parallel to the resultant of vectors 4i – 3j and 8i + 8j and 8i + 8j will be: (a) (24i+5j)/13 (b) (12i+5j)/13 (c) (6i+5j)/13 (d) none of these 3] A force of 6 N and another of 8 N can be applied together to produce the effect of a single force of: (a) 1 N (b) 11 N (c) 15 N (d) 20 N 4] Two vectors having equal magnitudes A make an angle ? with each other. The magnitude and direction of the resultant are respectively: (a) 2Acos ? /2, along bisector (b) A cos ?/2, at 45 from one vector (c) A cos ?/2, along bisector (d) A sin ?/2, along bisector


1]If A = B+C and the magnitude of A, B and C are 5,4,3 unit, the angle between A and B is :


(a)cos-13/5                          (b) cos-14/5


(c)π/2                                    (d) sin-1 3/5



2] Unit vector parallel to the resultant of vectors 4i – 3j and 8i + 8j and 8i + 8j will be:


(a) (24i+5j)/13                   (b) (12i+5j)/13


(c) (6i+5j)/13                      (d) none of these


3] A force of 6 N and another of 8 N can be applied together to produce the effect of a single force of:


(a) 1 N                   (b) 11 N


(c) 15 N                 (d) 20 N


4] Two vectors having equal magnitudes A make an angle ? with each other. The magnitude and direction of the resultant are respectively:


(a) 2Acos ? /2, along bisector


(b) A cos ?/2, at 45 from one vector


(c) A cos ?/2, along bisector


(d) A sin ?/2, along bisector

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Anand

1. A-B=C square both sides

25+ 16-40 cos@=9  , cos@=32/40 = 4/5 ans b

2. resultant vector sum of all three = 20i + 13j  hence d

3. resultant square= 36 + 64 + 96 cos@= 100 + 96 cos@ range 4 to 196 

 resultant 2 to 16 answer b

4. R2= 2A2 + 2A2 COS@= 2A2 (1+COS@)=2A2 2COS2@/2

     R= 2ACOS@/2 option a

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..